OpenStudy (anonymous):

which number is half way between 3/16 and 1 and 7/8 A=3/4 B=0 C=1 and 1/32 D= 1 and 1/8

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):

Making a sketch may help :)

OpenStudy (anonymous):

i think its D

OpenStudy (anonymous):

im sure im wrong though

OpenStudy (skullpatrol):

half-way between $$ \frac{3}{16}$$ and $$1\frac{7}{8}$$

OpenStudy (anonymous):

ok that makes more sense

OpenStudy (skullpatrol):

change the second mixed number into a fraction please

OpenStudy (skullpatrol):

$$1\frac{7}{8}=\frac{15}{8}$$

OpenStudy (anonymous):

wouldn't it be 15/8 ??

OpenStudy (anonymous):

and you would subtract right...

OpenStudy (skullpatrol):

now change the improper fraction 15/8 into a fraction with a denominator of 16 just like the first fraction

OpenStudy (skullpatrol):

$$\frac{15}{8}=\frac{15*2}{8*2}=\frac{30}{16}$$

OpenStudy (skullpatrol):

@Ryan5757 can you now find which number is half way between 3/16 and 30/16?

OpenStudy (skullpatrol):

$$\Huge\frac{\frac{3}{16}+\frac{30}{16}}{2}$$

OpenStudy (skullpatrol):

recall:$$ \frac{a}{b}=a*\frac{1}{b}$$

OpenStudy (skullpatrol):

@Ryan5757 Are you still there?

OpenStudy (anonymous):

yes i am writing this down sorry i am slow

OpenStudy (skullpatrol):

np :)

OpenStudy (anonymous):

so NOW i know its not D ,so i can eliminate that

OpenStudy (skullpatrol):

$$\Huge\frac{\frac{3}{16}+\frac{30}{16}}{2}=(\frac{3}{16}+\frac{30}{16})*\frac{1}{2}$$

OpenStudy (skullpatrol):

By the definition of division

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

sorry i am taking so long i am working on work i only have 8-13 assignments done and i need to have 18 by january 1st and im behind where my dad just died

OpenStudy (anonymous):

im just so far behind , lol i have not slept in 2 days lol

OpenStudy (skullpatrol):

get some sleep and finish this one when you wake up

OpenStudy (anonymous):

lol i can try but alot of stress haha

OpenStudy (skullpatrol):

try

OpenStudy (anonymous):

thanks ,i will thanks for ur help

OpenStudy (skullpatrol):

thanks for asking :)

OpenStudy (anonymous):

.